# Optimized Gradient Clipping for Noisy Label Learning

### v1

@@ Abstract
Previous research has shown that constraining the gradient of loss function w.r.t. model-predicted probabilities can enhance the model robustness against noisy labels. These methods typically specify a fixed optimal threshold for gradient clipping through validation data to obtain the desired robustness against noise. However, this common practice overlooks the dynamic distribution of gradients from both clean and noisy-labeled samples at different stages of training, significantly limiting the model capability to adapt to the variable nature of gradients throughout the training process. To address this issue, we propose a simple yet effective approach called Optimized Gradient Clipping (OGC), which dynamically adjusts the clipping threshold based on the ratio of noise gradients to clean gradients after clipping, estimated by modeling the distributions of clean and noisy samples. This approach allows us to modify the clipping threshold at each training step, effectively controlling the influence of noise gradients. Additionally, we provide statistical analysis to certify the noise-tolerance ability of OGC. Our extensive experiments across various types of label noise, including symmetric, asymmetric, instance-dependent, and real-world noise, demonstrate the effectiveness of our approach. The source code of our method is available at: https://github.com/Virusdoll/Optimized-Gradient-Clipping.
@@ 2 Related Work
Other methods for noisy label learning. In addition to robust loss functions, various other approaches are utilized for learning with noisy labels. A sample selection-based approach aims to identify and select clean samples for training, often leveraging the disagreement between two models (Malach and Shalev-Shwartz 2017; Jiang et al. 2018; Han et al. 2018b; Yu et al. 2019; Wei et al. 2020). Some research further combines these methods with techniques from unsupervised or semi-supervised learning to make full use of the unselected corrupted samples (Nguyen et al. 2020; Li, Socher, and Hoi 2020; Berthelot et al. 2019). A loss correction-based approach modifies the loss of each sample based on label-dependent weights (Natarajan et al. 2013), an estimated noise transition matrix (Han et al. 2018a; Patrini et al. 2017), or model predictions (Arazo et al. 2019). Additionally, a regularization-based approach is also employed to combat label noise, including methods such as MixUp (Zhang et al. 2018), label smoothing (Lukasik et al. 2020), or leveraging previous prediction results (Liu et al. 2020).
@@ 3.2 Gradient Clipping
$\qquad\bar{g}_{\theta}=\text{clip}(\bar{g}_{\theta},\tau),$
$\text{where}\quad\bar{g}_{\theta}=\frac{1}{M}\sum_{m=1}^{M}\nabla_{\theta}\ell(f(x_{m}),y_{m}).$ (4)
@@ 4.2 Method
Here, we formally introduce our method, named as Optimized Gradient Clipping (OGC). Our approach consists of the following three components: 1) Modeling clean and noisy distributions, 2) Determining the clipping threshold, and 3) Applying clipping to the loss functions. Our algorithm is outlined in Appendix B.1.
@@
$\min_{\tau^{(t)}}\tilde{r}^{(t)}(\tau^{(t)})\quad\text{s.t.}\quad\tilde{r}^{(t)}\geq 1+\epsilon$ (12)
@@
Table 1: Test accuracies (%) of different $\tau^{(t)}$ adjusting methods on CIFAR-10 dataset with different label noise. Bold denotes the best results.
@@ 4.3 Robustness Analysis
The proof is deffered to Appendix C.1. Proposition 1 demonstrates that Cross Entropy equipped with our proposed OGC is always bounded. Moreover, as $\tau^{(t)}$ approaches 1, the bound becomes tighter. When $\tau^{(t)}=1$, our loss function is equivalent to the symmetric Mean Absolute Error (MAE). Notably, this symmetry is a key property that makes a loss function robust to label noise (Ghosh, Kumar, and Sastry 2017).
Proposition 1 paves the way for further noise robustness analysis of CE+OGC loss ¯$\bar{\ell}_{\text{CE}}$. Firstly, under the common instance-independent setting, we can depict the bound of excess risk for both symmetric and asymmetric label noise (see Equation 1 for the definitions).
@@
Table 3: Test accuracies (%) of different methods on CIFAR-10 datasets with different label noise. The results (mean±$\pm$±std) are reported over 3 random runs under different random seeds (1, 2, 3). Bold denotes the best results and underline denotes the second-best results.
@@
Table 4: Test accuracies (%) of different methods on CIFAR-100 datasets with different label noise. The results (mean±$\pm$±std) are reported over 3 random runs under different random seeds (1, 2, 3). Bold denotes the best results and underline denotes the second-best results..
@@
The proofs for Theorem 1 and Theorem 2 can be found in Appendix C.2 and Appendix C.3, respectively. These theorems illustrate that, under both symmetric and asymmetric label noise, with our proposed OGC method, the excess risk between $\tilde{f}^{\star}$ and $f^{\star}$ is consistently bounded. Meanwhile, these bounds get tighter as $\tau^{(t)}$ decreases.
@@
The proof for Theorem 3 is provided in Appendix C.4. This theorem demonstrates that, even under instance-dependent label noise, our proposed OGC method ensures that the risk differences between $\tilde{f}^{\star}$ and $f^{\star}$ remain consistently bounded.
@@
Table 5: Top-1 validation accuracy (%) on the clean ILSVRC12 validation set and WebVision validation set of ResNet-18 models trained on WebVision using different loss functions, under the “Mini” setting (Jiang et al. 2018). Bold denotes the best results and underline denotes the second-best results.
@@ 5.1 Empirical Understanding
In this subsection, to enhance our understanding of the proposed OGC, we conduct a series of experiments. Unless otherwise stated, our experimental settings follow those described in Section 5.2. For additional experimental results on parameter analysis, including predefined gradient ratio $1+\epsilon$, queue $Q$, time frame $s$, and time consumption, please refer to Appendix A.1.
Comparison with manually designed methods. To highlight the advantages of our optimize-based strategy, we conduct a comparison on CIFAR-10 dataset against the following manually designed approaches for tuning $\tau^{(t)}$: 1) Linear Decrease Strategy: This approach is defined by $\tau^{(t)}=\beta\cdot(1-\frac{t}{T})$, where $T$ represents the total number of time steps and $\beta$ is a parameter. 2) Exponential Moving Average (EMA) Strategy: This approach uses the formula $\frac{1}{\tau^{(t)}}=\alpha\cdot\frac{1}{\tau^{(t-1)}}+(1-\alpha)\cdot 1$, where $\alpha$ is a parameter. We also compare our optimize-based strategy with a fixed strategy proposed by Menon et al. (2020), where $\tau^{(t)}$ remains constant.
The experimental results are presented in Table 1, and the experimental details can be found in Appendix A.1.
@@
Can OGC improves robust loss functions? Our proposed OGC can be applied not only to non-robust loss functions, such as CE and FL, but also to robust loss functions. This naturally raises the question of whether an existing robust loss function can improve its performance by incorporating our proposed OGC. Here, we simply take Generalized Cross Entropy (GCE) (Zhang and Sabuncu 2018) as an example and demonstrate that equipping with OGC can enhance the performance of a robust loss function.
@@ 5.2 Evaluation on Benchmark Datasets
In this subsection, we evaluate our proposed method using two well-known datasets: CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009). To demonstrate the robustness of our proposed method against different types of label noise, we compare it with other state-of-the-art methods under various label noise conditions, including symmetric, asymmetric, instance-dependent (Xia et al. 2020), and real-world (Wei et al. 2022) label noise settings.
@@
Experimental details. Noise generation, training, and parameter settings are in Appendix A.2.
Results. The results for CIFAR-10 and CIFAR-100 are presented in Table 3 and Table 4, respectively. The reported values reflect the average accuracies over the last 10 epochs. For both datasets, the incorporation of OGC enables CE and FL to consistently achieve the top-2 average accuracy across various scenarios. Notably, CE and FL are among the least robust loss functions, emphasizing OGC’s effectiveness in enhancing the performance of non-robust loss functions. Notably, CE and FL, typically among the least robust loss functions, show significant improvement with OGC, highlighting its effectiveness in enhancing the performance of non-robust loss functions. For instance, in Table 3, on CIFAR-10 with real-world label noise, integrating OGC with CE achieves a direct improvement (+17.02%). Moreover, in Table 4, with the integration of OGC, CE and FL consistently achieve the top two accuracies in all scenarios except for CIFAR-10 with asymmetric label noise. Additionally, due to our optimize-based strategy, our loss function requires only a single hyper-parameter, significantly reducing the time needed for hyper-parameter tuning compared to the previous SOTA method, NCE+NNCE, which involves three hyper-parameters. These observations underscore the effectiveness of the proposed OGC in mitigating the impact of label noise
@@ 5.3 Evaluation on Real-world Noisy Dataset
To evaluate the performance of our proposed method on large-scale a real-world noisy dataset, we conduct experiments on WebVision 1.0 dataset. WebVision 1.0 (Li et al. 2017) contains more than 2.4 million web images crawled from the internet by using queries generated from the 1,000 class labels of the ILSVRC 2012 (Deng et al. 2009) benchmark. Following the “Mini” setting in previous works (Jiang et al. 2018), we only take the first 50 classes of the Google resized image subset. We evaluate the trained networks on the same 50 classes of both the ILSVRC 2012 validation set and WebVision 1.0 validation set, these can be considered as clean validation sets.
The experimental details can be found in Appendix A.3. The results of last epoch validation accuracies are reported in Table 5. As observed, integrating our proposed OGC with CE and FL results in significant improvements, outperforming existing robust loss functions on both validation sets. These results verify that our method is effective in enhancing noise-robustness in large-scale real-world scenarios.
@@ 6 Conclusions
In this paper, we introduce Optimized Gradient Clipping (OGC), a novel method for noisy label learning, which employs a 2-component Gaussian Mixture Model (2-GMM) to distinguish between clean and noisy samples, estimating the ratio of noise gradients to clean gradients after clipping. This approach allows us to set a clipping threshold at each training step, restricting the gradient ratio to a predefined limit and effectively limiting the influence of noise gradients. As a result, OGC enhances the robustness of both non-robust and robust loss functions when dealing with noisy labels. We also provide a theoretical analysis highlighting the noise tolerance of our proposed OGC. Extensive experiments demonstrate the effectiveness of OGC across various types of noise.
7 Acknowledgments
This work is supported by the Shanghai Engineering Research Center of Intelligent Computing System (No. 19DZ2252600) and the Research Grants Council (RGC) under grant ECS-22303424.
Appendix A Experiments
A.1 Empirical Understanding
@@
A.2 Evaluation on Benchmark Datasets
Noise generation. For class-independent noise, the noisy labels are generated following standard approaches in previous works (Ma et al. 2020; Zhou et al. 2021). For symmetric noise, we flip the labels in each class randomly to incorrect labels of other classes. For asymmetric noise, we flip the labels within a specific set of classes. For CIFAR-10, flipping TRUCK $\to$ AUTOMOBILE, BIRD $\to$ AIRPLANE, DEER $\to$ HORSE, CAT $\leftrightarrow$ DOG. For CIFAR-100, the 100 classes are grouped into 20 super-classes with each has 5 sub-classes, and each class are flipped within the same super-class into the next in a circular fashion. We vary the noise rate $\eta\in\{0.2,0.5,0.8\}$ for symmetric noise and $\eta\in\{0.4\}$ for asymmetric noise. For instance-dependent noise, we use the part-dependent noise from PDN (Xia et al. 2020) with a noise rate of 40%, where the noise is synthesized based on the DNN prediction error. For real-world noise, we use the “Worst” label set of CIFAR-10N and the “Noisy” label set of CIFAR-100N (Wei et al. 2022), respectively.
@@ A.2 Evaluation on Benchmark Datasets
A.3 Evaluation on Real-world Noisy Dataset
@@
Appendix B Method
B.1 Algorithm
Our algorithm is presented in Algorithm 1. For an in-depth discussion on the queue $Q$ and the time frame $s$, please refer to Appendix A.1.
Algorithm 1 Training process for Optimized Gradient Clipping (OGC)
2:for   do
4:     for $m=1$ to $M$ do
6:         Enqueue($Q,H_{m}$) $\triangleright$ Add cross entropy value to queue
7:     end for
8:     if $t\mod s=0$ then
[…]
11:     else
13:     end if
16:end for
17:return Trained model $f$
B.2 Limitations

### v4

@@ Abstract
Previous research has shown that constraining the gradient of loss function w.r.t. model-predicted probabilities can enhance the model robustness against noisy labels. These methods typically specify a fixed optimal threshold for gradient clipping through validation data to obtain the desired robustness against noise. However, this common practice overlooks the dynamic distribution of gradients from both clean and noisy-labeled samples at different stages of training, significantly limiting the model capability to adapt to the variable nature of gradients throughout the training process. To address this issue, we propose a simple yet effective approach called Optimized Gradient Clipping (OGC), which dynamically adjusts the clipping threshold based on the ratio of noise gradients to clean gradients after clipping, estimated by modeling the distributions of clean and noisy samples. This approach allows us to modify the clipping threshold at each training step, effectively controlling the influence of noise gradients. Additionally, we provide statistical analysis to certify the noise-tolerance ability of OGC. Our extensive experiments across various types of label noise, including symmetric, asymmetric, instance-dependent, and real-world noise, demonstrate the effectiveness of our approach.
Codehttps://github.com/Virusdoll/OGC
@@ 2 Related Works
Other methods for noisy label learning. In addition to robust loss functions, various other approaches are utilized for learning with noisy labels. A sample selection-based approach aims to identify and select clean samples for training, often leveraging the disagreement between two models (Yu et al. 2019; Wei et al. 2020). Some research further combines these methods with techniques from unsupervised or semi-supervised learning to make full use of the unselected corrupted samples (Nguyen et al. 2020; Li, Socher, and Hoi 2020; Wei et al. 2022b). A loss correction-based approach modifies the loss of each sample based on label-dependent weights (Natarajan et al. 2013), an estimated noise transition matrix (Han et al. 2018; Patrini et al. 2017), or model predictions (Arazo et al. 2019). Additionally, a regularization-based approach is also employed to combat label noise, including methods such as MixUp (Zhang et al. 2018), label smoothing (Lukasik et al. 2020), or leveraging previous prediction results (Liu et al. 2020).
@@ 3.2 Gradient Clipping
$\qquad\bar{g}_{\theta}=\texttt{clip}(g_{\theta},\tau),$
$\text{where}\quad g_{\theta}=\frac{1}{M}\sum_{m=1}^{M}\nabla_{\theta}\ell(f(x_{m}),y_{m}).$ (4)
@@ 4.2 Method
Here, we formally introduce our method, named as Optimized Gradient Clipping (OGC). Our approach consists of the following three components: 1) Modeling clean and noisy distributions, 2) Determining the clipping threshold, and 3) Applying clipping to the loss functions. Our algorithm is outlined in Appendix.
@@
$\min_{\tau^{(t)}}\tilde{r}^{(t)}(\tau^{(t)})\quad\text{s.t.}\quad\tilde{r}^{(t)}\geq 1+\epsilon,$ (12)
@@
Table 1: Test accuracies (%) of $\tau^{(t)}$ adjusting methods on CIFAR-10 dataset with different label noise, bold for best results.
@@ 4.3 Robustness Analysis
The proof is deffered to Appendix. Proposition 1 demonstrates that Cross Entropy equipped with our proposed OGC is always bounded. Moreover, as $\tau^{(t)}$ approaches 1, the bound becomes tighter. When $\tau^{(t)}=1$, our loss function is equivalent to the symmetric Mean Absolute Error (MAE). Notably, this symmetry is a key property that makes a loss function robust to label noise (Ghosh, Kumar, and Sastry 2017).
Proposition 1 paves the way for further noise robustness analysis of CE+OGC loss ¯$\bar{\ell}_{\text{CE}}$. Firstly, under the common instance-independent setting, we can depict the bound of excess risk for both symmetric and asymmetric label noise (see Equation (1) for the definitions).
@@
Table 3: Test accuracies (%) of different methods on CIFAR-10 datasets with different label noise. The results (mean±$\pm$±std) are reported over 3 random runs. Bold denotes the best results and underline denotes the second-best results.
@@
Table 4: Test accuracies (%) of different methods on CIFAR-100 datasets with different label noise. The results (mean±$\pm$±std) are reported over 3 random runs. Bold denotes the best results and underline denotes the second-best results.
@@
The proofs for Theorem 1 and Theorem 2 can be found in Appendix. These theorems illustrate that, under both symmetric and asymmetric label noise, with our proposed OGC method, the excess risk between $\tilde{f}^{\star}$ and $f^{\star}$ is consistently bounded. Meanwhile, these bounds get tighter as $\tau^{(t)}$ decreases.
@@
The proof for Theorem 3 is provided in Appendix. This theorem demonstrates that, even under instance-dependent label noise, our proposed OGC method ensures that the risk differences between $\tilde{f}^{\star}$ and $f^{\star}$ remain consistently bounded.
@@
Table 5: Top-1 validation accuracies (%) of different methods on the ILSVRC12 and WebVision validation sets, under the “Mini” setting (Jiang et al. 2018). Bold denotes the best results and underline denotes the second-best results.
@@ 5.1 Empirical Understanding
In this subsection, to enhance our understanding of the proposed OGC, we conduct a series of experiments. Unless otherwise stated, our experimental settings follow those described in Section 5.2. For additional experimental results on parameter analysis, including predefined gradient ratio $1+\epsilon$, queue $Q$, time frame $s$, and time consumption, please refer to Appendix.
Comparison with manually designed methods. To highlight the advantages of our optimize-based strategy, we compare it on CIFAR-10 against manually designed approaches for tuning $\tau^{(t)}$: 1) Linear Decrease Strategy: This approach is defined by $\tau^{(t)}=\beta\cdot(1-\frac{t}{T})$, where $T$ represents the total number of time steps and $\beta$ is a parameter. 2) Exponential Moving Average (EMA) Strategy: This approach uses the formula $\frac{1}{\tau^{(t)}}=\alpha\cdot\frac{1}{\tau^{(t-1)}}+(1-\alpha)\cdot 1$, where $\alpha$ is a parameter. We also compare our optimize-based strategy with a fixed strategy proposed by Menon et al. (2020), where $\tau^{(t)}$ remains constant.
The experimental results are presented in Table 1, and the experimental details can be found in Appendix.
@@
Can OGC improves robust loss functions? Our proposed OGC can be applied not only to non-robust loss functions, such as CE and FL, but also to robust loss functions. This naturally raises the question of whether an existing robust loss function can improve its performance by incorporating our proposed OGC. Here, we use Generalized Cross Entropy (GCE) (Zhang and Sabuncu 2018) as an example and demonstrate that equipping with OGC can enhance the performance of a robust loss function.
@@ 5.2 Evaluation on Benchmark Datasets
In this subsection, we evaluate our proposed method using two well-known datasets: CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009). To demonstrate the robustness of our proposed method against different types of label noise, we compare it with other state-of-the-art methods under various label noise conditions, including symmetric, asymmetric, instance-dependent (Xia et al. 2020), and real-world (Wei et al. 2022a) label noise settings.
@@
Experimental details. Noise generation, training, and parameter settings are in Appendix.
Results. The results for CIFAR-10 and CIFAR-100 are presented in Table 3 and Table 4, respectively. The reported values reflect the average accuracies over the last 10 epochs. For both datasets, the incorporation of OGC enables CE and FL to consistently achieve the top-2 average accuracy across various scenarios. Notably, CE and FL are among the least robust loss functions, emphasizing OGC’s effectiveness in enhancing the performance of non-robust loss functions. Notably, CE and FL, typically among the least robust loss functions, show significant improvement with OGC, highlighting its effectiveness in enhancing the performance of non-robust loss functions. For instance, in Table 3, on CIFAR-10 with real-world label noise, integrating OGC with CE achieves a direct improvement (+17.02%). Moreover, in Table 4, with the integration of OGC, CE and FL consistently achieve the top two accuracies in all scenarios except for CIFAR-10 with asymmetric label noise. Additionally, due to our optimize-based strategy, our loss function requires only a single hyper-parameter, significantly reducing the time needed for hyper-parameter tuning compared to the previous SOTA method, NCE+NNCE, which involves three hyper-parameters. These observations underscore the effectiveness of our proposed OGC.
@@ 5.3 Evaluation on Real-world Noisy Dataset
To evaluate our proposed method on large-scale real-world noisy dataset, we conduct experiments on WebVision 1.0 dataset. WebVision 1.0 (Li et al. 2017) contains more than 2.4 million web images crawled from the internet by using queries generated from the 1,000 class labels of the ILSVRC 2012 (Deng et al. 2009) benchmark. Following the “Mini” setting in previous works (Jiang et al. 2018), we only take the first 50 classes of the Google resized image subset. We evaluate the trained networks on the same 50 classes of both the ILSVRC 2012 validation set and WebVision 1.0 validation set, these can be considered as clean validation sets.
The experimental details can be found in Appendix. The results of last epoch validation accuracies are reported in Table 5. As observed, integrating our proposed OGC with CE and FL results in significant improvements, outperforming existing robust loss functions on both validation sets. These results verify that our method is effective in enhancing noise-robustness in large-scale real-world scenarios.
@@ 6 Conclusions
In this paper, we introduce Optimized Gradient Clipping (OGC), a novel method for noisy label learning, which employs a 2-component Gaussian Mixture Model (2-GMM) to distinguish between clean and noisy samples, estimating the noise-to-clean gradient ratio after clipping. This allows us to set a clipping threshold at each training step, limiting the influence of noise gradients. OGC enhances the robustness of both non-robust and robust loss functions in noisy label settings. Our theoretical analysis highlights its noise tolerance, and extensive experiments demonstrate its effectiveness across various noise types.
Acknowledgments
This work is supported by the Shanghai Engineering Research Center of Intelligent Computing System (grant No. 19DZ2252600) and the Research Grants Council (RGC) under grant ECS-22303424.
Appendix A Method
A.1 Algorithm
Our algorithm is presented in Algorithm 1. For an in-depth discussion on the queue $Q$ and the time frame $s$, please refer to Appendix B.1.
Algorithm 1 Training process for Optimized Gradient Clipping (OGC)
2:for $t=1$ to $T$ do
4:     for $m=1$ to $M$ do
6:         Enqueue($Q,H_{m}$) $\triangleright$ Add cross entropy value to queue
7:     end for
8:     if $t\mod s=0$ then
[…]
11:     else
13:     end if
16:end for
17:return Trained model $f$
A.2 Limitations
Appendix B Experiments
B.1 Empirical Understanding
@@
Can OGC improves robust loss functions? In addition to the discussion in main text, we provide the hyper-parameters used in Table 2. For $q$, we adopted the optimal value as reported in the GCE paper (Zhang and Sabuncu 2018). For $\epsilon_{0}$, we tuned it over $\{100,50,20,10\}$, following the tuning procedure outlined in Appendix B.2, and selected the value that performed best under 0.5 symmetric label noise.
@@
B.2 Evaluation on Benchmark Datasets
Noise generation. For class-independent noise, the noisy labels are generated following standard approaches in previous works (Ma et al. 2020; Zhou et al. 2021). For symmetric noise, we flip the labels in each class randomly to incorrect labels of other classes. For asymmetric noise, we flip the labels within a specific set of classes. For CIFAR-10, flipping TRUCK $\to$ AUTOMOBILE, BIRD $\to$ AIRPLANE, DEER $\to$ HORSE, CAT $\leftrightarrow$ DOG. For CIFAR-100, the 100 classes are grouped into 20 super-classes with each has 5 sub-classes, and each class are flipped within the same super-class into the next in a circular fashion. We vary the noise rate $\eta\in\{0.2,0.5,0.8\}$ for symmetric noise and $\eta\in\{0.4\}$ for asymmetric noise. For instance-dependent noise, we use the part-dependent noise from PDN (Xia et al. 2020) with a noise rate of 40%, where the noise is synthesized based on the DNN prediction error. For real-world noise, we use the “Worst” label set of CIFAR-10N and the “Noisy” label set of CIFAR-100N (Wei et al. 2022a), respectively.
@@ B.2 Evaluation on Benchmark Datasets
B.3 Evaluation on Real-world Noisy Dataset